Please show all the works. Thanks 7. Award: 10.00 points A pension fund manager is considering three mutual funds. The first is a stock fund, the second is a l
Answer #1
To find the fraction of wealth to invest in Stock fund that will result in the risky portfolio with maximum Sharpe ratio
the following formula to determine the weight of Stock fund in risky portfolio should be used
w(*d)= ((E[Rd]-Rf)*Var(Re)-(E[Re]-Rf)*Cov(Re,Rd))/((E[Rd]-Rf)*Var(Re)+(E[Re]-Rf)*Var(Rd)-(E[Rd]+E[Re]-2*Rf)*Cov(Re,Rd)
Where
Stock fund E[R(d)]= 12.00%
Bond fund E[R(e)]= 5.00%
Stock fund Stdev[R(d)]= 41.00%
Bond fund Stdev[R(e)]= 30.00%
Var[R(d)]= 0.16810
Var[R(e)]= 0.09000
T bill Rf= 3.00%
Correl Corr(Re,Rd)= 0.0667
Covar Cov(Re,Rd)= 0.0082
Stock fund Therefore W(*d)= 0.7515
Bond fund W(*e)=(1-W(*d))= 0.2485
Expected return of risky portfolio= 10.26%
Risky portfolio std dev (answer Risky portfolio std dev)= 32.18%
Where
Var = std dev^2
Covariance = Correlation* Std dev (r)*Std dev (d)
Expected return of the risky portfolio = E[R(d)]*W(*d)+E[R(e)]*W(*e)
Risky portfolio standard deviation =( w2A*σ2(RA)+w2B*σ2(RB)+2*(wA)*(wB)*Cor(RA,RB)*σ(RA)*σ(RB))^0.5
Desired return = tbill return*proportion invested in tbill+risky portfolio return *proportion invested in risky portfolio
= tbill return*proportion invested in tbill+risky portfolio return *(1-proportion invested in tbill)
0.09=0.03*Proportion invested in Tbill+0.1026*(1-Proportion invested in Tbill)
Proportion invested in Tbill (answer b-1) = (0.1026-0.09)/(0.1026-0.03)
=0.1736 (17.36%)
proportion invested in risky portfolio = 1-proportion invested in tbill
=0.8264 (82.64%)
Proportion invested in Bond fund (answer proportion invested in Bond fund) =proportion invested in risky portfolio *weight of Bond fund
=0.2053 (20.53%)
Proportion invested in Stock fund (answer b-2) =proportion invested in risky portfolio *weight of Stock fund
=0.6211 (62.11%)
std dev of portfolio (answer a) = std of risky portfolio*proportion invested in risky portfolio
0.8264*0.3218=26.59%